OpenStudy (anonymous):

domain of \[\sin^{-1} [2-4x^2]\] where [...] is greatest interger function

ganeshie8 (ganeshie8):

we need to solve \[-1 \le \lfloor 2-4x^2 \rfloor \le 1\]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\[-1 \le 2-4x^2 \lt 2\]

ganeshie8 (ganeshie8):

\[\frac{-\sqrt{3}}{2} \le x \lt \frac{\sqrt{3}}{2}\] \(x \ne 0\)

ganeshie8 (ganeshie8):

see if that looks okay ^